Area of parallelogram is 2019 square unite. Leanth and width of it is integer. then, what is the minimum perimeter value for it?
1
Expert's answer
2020-02-13T10:13:17-0500

Solution. The parallelogram area can be calculated using the formula


"S=aH_a=bH_b=ab\times sin\alpha"

where a, b are sides parallelogram; α is the angle between the sides of the parallelogram; Ha and Hb are parallelogram heights. The perimeter will be minimal if the sides of the parallelogram are perpendicular. Therefore


"S=ab"

By factoring the number 2019, we get


"2019=3\times673"

Numbers 3 and 673 are prime.

As result the minimum perimeter value  is equal to


"P=2(3+673)=1352"

Answer. 1352.